# Rendezvous in CAVITY: Kinematics and gas properties of an isolated dwarf-dwarf merging pair in a cosmic void region

### v1

@@ Abstract
Context. Galaxy mergers are pivotal events in the evolutionary history of galaxies, with their impact believed to be particularly significant in dwarf galaxies due to their low masses. However, these events remain largely underexplored, especially in pristine environments like voids.
@@
Methods. Using CAVITY PPAK-IFU data combined with deep optical broadband imaging from the Isaac Newton Telescope, we analysed the kinematics and ionized gas properties of each dwarf galaxy in the system by employing a full spectral fitting technique.
Results. The orientation of this merging pair relative to the line of sight allowed us to determine the dynamical mass of each component, showing that both had similar dynamical masses within galactocentric distances of up to 2.9 kpc. These galaxies were likely star-forming dwarfs with rotating discs prior to the merger. While the gas-phase metallicity of both components is consistent with that of star-forming dwarf galaxies, the star formation rates observed in both components exceed those typically reported for equally massive star-forming dwarf galaxies. This indicates that the merger has presumably contributed to enhancing star formation. Our analysis shows no signs of AGN activity in this merging pair. Furthermore, we found no significant difference in the optical g-r colour of this merging pair compared to other merging dwarf pairs across different environments.
Conclusions. While most merging events occur in group-like environments with high galaxy density and the tidal influence of a host halo, and isolated mergers typically involve galaxies with significant mass differences, the identified merging pair does not follow these patterns. We speculate that the global dynamics of the void or past three-body encounters involving components of this pair and a nearby dwarf galaxy might have triggered this merging event.
@@
Merger events play a crucial role in the mass growth and evolution of galaxies as they significantly impact their physical properties, including their star formation rate (SFR), morphology, and black hole growth (e.g., Toomre & Toomre, 1972; Hopkins et al., 2010). The hierarchical cold dark matter (CDM) paradigm (Cole et al., 2000) predicts that merger events are frequent throughout the evolutionary history of galaxies, occurring commonly among low-mass ones, with log(M⋆/M⊙) 9.5 (e.g., Klimentowski et al., 2010; Deason et al., 2014). Recent observations increasingly support the prevalence of such events among dwarf galaxies (e.g., Martínez-Delgado et al., 2012; Rich et al., 2012; Amorisco et al., 2014; Crnojević et al., 2014). Also, N-body simulations, such as Millennium-II, predict at least three major mergers111Merging between galaxies of comparable stellar masses. for dwarf galaxies during their lifetime (Fakhouri et al., 2010). Deason et al. (2014) showed that more than 10 percent of dwarf galaxies, as satellites of Milky Way (MW) or M31-like host halos, underwent at least one major merger event since redshift , and prior to infall to their present-day host halo or shortly ( 1-3 Gyr) after that (see also Wetzel et al., 2015). Beyond the virial radius of MW or M31-like host halos, and in more isolated environments, the anticipated number of recent merger events doubles. Martin et al. (2021) showed that isolated dwarf galaxies and those in galaxy groups had experienced at least one minor and one major merger during their lifetime, between redshift 0.5 z 5.0.
@@
The impact of merger events is expected to be particularly pronounced in dwarf galaxies due to their low mass. Minor and major mergers, along with internal processes and tidal interactions, are believed to be the primary drivers of the permanent morphological transformations in dwarfs (e.g., Toomre, 1977; Knebe et al., 2006; Dekel et al., 2009; Naab et al., 2014; Martin et al., 2018). Furthermore, mergers are believed to be the main trigger of intense star formation in pairs (e.g., Tissera et al., 2002; Ellison et al., 2013), some of which result in the formation of starburst dwarfs such as Blue Compact Dwarf galaxies (BCDs, Zhang et al., 2020). Similarly, Ruiz-Lara et al. (2021) interpreted the reignition of star formation after a quiescent period and the presence of metal-poor stars in Leo I as evidence of a past merger event, possibly with an ultra-faint dwarf galaxy. The dwarf-dwarf mergers can also leave discernible traces in their gas-phase metallicity. For instance, Michel-Dansac et al. (2008) showed that in a large sample of dwarf galaxy pairs with strong signs of interactions, the gas-phase metallicity is 0.2 dex higher than in their non-interacting counterparts.
@@
Observed morphological asymmetries and disturbances, such as shell-like features and tidal tails around dwarf galaxies, serve as a telltale of their relatively recent interactions (e.g., Conselice et al., 2003; Klimentowski et al., 2010; Paudel et al., 2017). For instance, Coleman et al. (2004) suggested a past merger as a possible source of the small shell-like feature around the Fornax dwarf galaxy (Yozin & Bekki, 2012). Amorisco et al. (2014) showed that the kinematically detected stellar stream in Andromeda II can be a possible remnant of a past dwarf-dwarf merger. Similarly, peculiarities in kinematics and chemical properties of other Local Group (LG) dwarf galaxies (such as Carina and Sculptor) are better explained as footprints of dwarf-dwarf encounters (Tolstoy et al., 2004; Venn et al., 2012).
@@
Detection of mergers among dwarfs extends beyond the LG (e.g., Paudel & Sengupta, 2017; Mićić et al., 2023; Román et al., 2023). For instance, Graham et al. (2012) proposed that the merger of two dwarf galaxies featuring disc-like substructures could account for the rectangular shape of LEDA074886. Paudel et al. (2017) investigated the merger origin of shell-like features in Virgo early-type dwarf galaxies. Subsequently, in Paudel et al. (2018), they shed light on the prevalence of dwarf-dwarf mergers by presenting an extensive catalogue of nearby interacting/merging dwarf galaxies up to z 0.02, but no information regarding the large-scale environment of this sample was given. Kimbro et al. (2021) highlighted the 10 kpc string of massive star clusters, with stellar populations younger than 10 Myr, bridging the northern and southern components of Mrk709 and suggested that it is possibly due to the early stages of a dwarf-dwarf galaxy merger.
@@
Most of the dwarf-dwarf mergers investigated in the literature thus far were mainly located in the high-density regions (i.e., in galaxy groups and outskirts of clusters), where environmental effects are vigorous and non negligible. In these cases, dwarf merging pairs are not only transformed due to the merging and close interaction with each other but also experience tidal forces exerted by their host halo as well as hydro-dynamical interactions with surrounding intergalactic medium (e.g., Boselli & Gavazzi, 2014). Therefore, in such cases, disentangling between other environmental transformations and those solely induced by the merger event is not straightforward. One way to deepen our comprehension of dwarf galaxies growth and transformation, specifically through mergers, is to investigate dwarf-dwarf mergers in isolation (e.g., Kado-Fong et al., 2020; Grajales-Medina et al., 2023; Calderón-Castillo et al., 2024; Calderón-Castillo & Smith, 2024; Vásquez-Bustos et al., 2025) and in pristine environments, such as cosmic voids (Annibali et al., 2016). Voids represent the most under-dense regions of the cosmic web (e.g., Peebles, 2001; Kreckel et al., 2011) spanning large volumes (i.e., the average size of 35 h-1 Mpc, van de Weygaert & Platen, 2011). Within these vast expanses, galaxies are less disturbed and modified by the complex environmental processes present in the galaxy groups and clusters.
@@
Few theoretical studies have focused on galaxy mergers in voids, revealing that the mean number of mergers does not depend significantly on the void environment (Rosas-Guevara et al., 2022). The only difference reported so far between mergers in voids and those in clusters or groups is that mergers in voids tend to occur more recently (i.e., 2 Gyr), particularly for galaxies with log(M⋆/M⊙) ¡ 9.5 (Rodríguez-Medrano et al., 2024). From observations, several studies also report the presence of dwarf galaxies in voids with unusual features in their morphologies, that could be indicative of past mergers (e.g., Kreckel et al., 2011). Interactions and mergers between small groups of galaxies are also reported in some voids by a few studies (e.g., Beygu et al., 2013).
@@
In this study, we report the serendipitous detection of a gas-rich merging dwarf pair located at the central regions (i.e., at R 0.13 of void effective radius) of a cosmic void region that has been observed as part of the Calar Alto Void Integral-field Treasury Survey (CAVITY222https://cavity.caha.es/) project (Pérez et al., 2024). This legacy integral-field unit (IFU) survey aims to investigate the formation and evolution of nearby (0.005 z 0.050) void galaxies using their spatially resolved kinematics and stellar population properties (e.g., Domínguez-Gómez et al., 2023; Conrado et al., 2024). The stellar mass ratio of galaxies in this merging pair, their local and large-scale environments, and their orientation with respect to the line of sight make it a great candidate for better understanding details of merger events in low-mass galaxies. To this end, we investigate spatially resolved (on kpc scales) ionised gas kinematics, gas-phase metallicity, and SFR of this isolated merging pair and compare our results with their other counterparts in denser environments than voids. The results of this study can shed light on the crucial role of mergers in shaping the present-day properties of dwarf galaxies and better distinguish between footprints of environmental effects and those stemming from merger events in the mass assembly of dwarf galaxies.
@@
This paper is organized as follows: in Section 2, we introduce the IFU and deep photometric data sets used for the analysis. In Section 3, we describe the methods we utilised for deriving kinematics and ionized gas emission. In Section 4, we explain the results, and in Section 5, we discuss them by comparing the interacting pair with its other counterparts from different environments in the literature. We summarise the main key points of this work in Section 6. In this work we assume a flat CDM cosmology with H0=69.6 [km s-1 Mpc-1], = 0.286, and = 0.714.
@@ 2.1 IFU data
The interacting pair was observed for a total exposure time of 5400 s divided into three dithering positions, each with two exposures of 900 s, using the PMAS/PPAK-IFU spectrograph at the 3.5 meter telescope of the Calar Alto Observatory (CAHA) as part of the CAVITY project and under the survey ID of CAVITY35844 (Pérez et al., 2024). Observations were carried out using the V500 grating of the IFU instrument, which covers the optical wavelength range from 3750 to 7500 Å with a resolving power of R 850 at 5000 Å and an average full width at half maximum (FWHM) of 6.0 Å. The instrument has a field-of-view (FOV) of 74 64 arcsec2, with a spatial sampling of 1 arcsec, which provides adequate coverage of both interacting pairs and their tidal tail (see Fig. 1). Due to the small size of the galaxy pair, it occupies only the central region of the PMAS FOV and is therefore unaffected by vignetting effects. Data reduction and construction of the final calibrated data cube were carried out through the CAVITY standard data reduction pipeline. In this process, along with other corrections and calibrations, the datacube has also been corrected for the Galactic extinction. For more details on the CAVITY data reduction pipeline, see García-Benito et al. (2024).
@@ 2.2 CAVITY+ deep optical imaging
For our analysis, we take advantage of the deep optical imaging carried out with the 2.54 m Isaac Newton Telescope (INT), as part of the CAVITY+ survey (Pérez et al., 2024). This survey uses the Wide Field Camera (WFC), mounted on the INT primary focus, which approximately covers a FOV of 34’ 34’ with a pixel scale of 0.33 arcsec pix-1. In Fig. 1 Panel A, we show CAVITY35843 and 35844 (marked with an orange box) along with CAVITY38680 (their only neighbouring galaxy within a projected search area of 1.5 Mpc, marked with a green box, see Table 1) in the reduced and calibrated INT images taken in the r-band filter. For details on the CAVITY+ deep optical imaging survey and data reduction, see Pérez et al. (2024).
@@
In Panel C of Fig. 1, we show a zoom-in view of CAVITY35843 and 35844 from DECaLS (left-hand Panel) and INT g+r images (right-hand Panel). While the data coming from INT have much higher surface brightnesses (30.0 and 29.0 mag arcsec-2 in g and r bands respectively, measured in 3 sigma, 1010 arcsec boxes; Román et al., 2020), their seeing is mediocre (1.9 and 2.0 arcsec FWHM). To address this, we utilize DECAM instrument data333Dark Energy Camera, https://noirlab.edu/science/programs/ctio/instruments/Dark-Energy-Camera, which have slightly lower surface brightness limits (approximately 1 mag arcsec-2 less Martínez-Delgado et al. 2023), but offer superior seeing conditions, with FWHM values of 1.18 and 1.29 arcsec in the r and g bands, respectively. Both data show two bright cores, representing the core of each interacting dwarf component, and a faint tidal tail heading toward the north-west. In Panel D of the same figure, we show the surface brightness map of this interacting pair using g-band INT images. The two bright cores of this interacting pair have similar surface brightness 19.5 [mag arcsec 20.5 in g-band. The tidal tail is 4 to 5 mag fainter than the two cores. Additionally, we could trace two low surface brightness (LSB) features ( 25 mag arcsec-2, marked with arrows in panel D) in the north-east and south-west of the interacting system, which are possibly the results of the ongoing interaction between these two dwarf galaxies. These LSB features are not traced in any other optical data available in the literature (including DECaLS presented here).
@@
In Panel E of Fig. 1, we show the g-r colour map of CAVITY35843 and 35844, computed using INT deep images in SDSS g and r filters. Blue regions with 0.10 g-r[mag]0.25, possibly dominated by star formation, are seen in the tidal tail and the galaxy’s east and west parts. Additionally, we trace a curved strip-like feature, dominated by red colours (g-r 0.6 mag), that separates the two bright cores of the interacting pair and their tidal tails. Compared with the DECaLS colour image in Panel C, one can notice that this red curved strip traces the distribution of dust. In the southern part of the interacting pair, we also observe a distribution of redder light. However, no dust lanes were visually detectable in these regions in the DECaLS data, hence it is possible that they trace presence of older and/or metal-richer stellar populations.
@@
In panels A and B of Fig. 2, we show the surface brightness (from the INT g-band image) and signal-to-noise ratio (S/N, extracted from PPAK data) maps of the merging pair, respectively. Here, we convolved the INT surface brightness map to match the spatial resolution of the PPAK data (i.e., with a pixel size of 1 arcsec corresponding to the spatial resolution of 2.5 arcsec). The S/N is defined on the continuum level of each spaxel’s spectrum and within the 4590 to 4610 Å wavelength range. As marked on the surface brightness map and to ease the discussion hereafter, we refer to the brightest interacting galaxy (i.e., CAVITY35843) as component A; the other (i.e., CAVITY35844) will be referred to as component B. In all of our analysis we excluded spaxels of the PPAK data with S/N 15 (except for analysis in Section 4.1.1) to attain reliable measurements. This mainly excludes the outskirts and faint tidal tail of the merging pair.

### v2

@@ Abstract
Context. Galaxy mergers are pivotal events in the evolutionary history of galaxies, with their impact believed to be particularly significant in dwarf galaxies due to their low masses. However, these events remain largely underexplored, especially in pristine environments such as voids.
@@
Methods. Using CAVITY PPAK-IFU data combined with deep optical broadband imaging from the Isaac Newton Telescope, we analysed the kinematics and ionised gas properties of each dwarf galaxy in the system by employing a full spectral fitting technique.
Results. The orientation of this merging pair relative to the line of sight allowed us to determine the dynamical mass of each component, which we found to have similar dynamical masses within galactocentric distances of up to 2.9 kpc. These galaxies were likely star-forming dwarfs with rotating discs prior to the merger. While the gas-phase metallicity of both components is consistent with that of star-forming dwarf galaxies, the star formation rates observed in both components exceed those typically reported for equally massive star-forming dwarf galaxies. This indicates that the merger has presumably contributed to enhancing star formation. Our analysis shows no signs of AGN activity in this merging pair. Furthermore, we found no significant difference in the optical g-r colour of this merging pair compared to other merging dwarf pairs across different environments.
Conclusions. While most merging events occur in group-like environments with a high galaxy density and the tidal influence of a host halo, and isolated mergers typically involve galaxies with significant mass differences, the identified merging pair does not follow these patterns. We speculate that the global dynamics of the void or past three-body encounters involving components of this pair and a nearby dwarf galaxy might have triggered this merging event.
@@
Merger events play a crucial role in the mass growth and evolution of galaxies, as they significantly impact their physical properties, including their star formation rate (SFR), morphology, and black hole growth (e.g. Toomre & Toomre, 1972; Hopkins et al., 2010). The hierarchical cold dark matter (CDM) paradigm (Cole et al., 2000) predicts that merger events are frequent throughout the evolutionary history of galaxies, occurring commonly among low-mass ones with log(M⋆/M⊙) 9.5 (e.g. Klimentowski et al., 2010; Deason et al., 2014). Recent observations increasingly support the prevalence of such events among dwarf galaxies (e.g. Martínez-Delgado et al., 2012; Rich et al., 2012; Amorisco et al., 2014; Crnojević et al., 2014). Also, N-body simulations, such as Millennium-II, predict at least three major mergers111Major mergers are mergers between galaxies of comparable stellar masses. for dwarf galaxies during their lifetime (Fakhouri et al., 2010). Deason et al. (2014) shows that more than 10 of dwarf galaxies, such as satellites of the Milky Way (MW) or M31-like host halos, underwent at least one major merger event since redshift , and prior to infall to their present-day host halo or shortly ( 1-3 Gyr) after that (see also Wetzel et al., 2015). Beyond the virial radius of the MW or M31-like host halos, and in more isolated environments, the anticipated number of recent merger events doubles. Martin et al. (2021) shows that isolated dwarf galaxies and those in galaxy groups had experienced at least one minor and one major merger during their lifetime between redshift 0.5 z 5.0.
@@
The impact of merger events is expected to be particularly pronounced in dwarf galaxies due to their low mass. Minor and major mergers, along with internal processes and tidal interactions, are believed to be the primary drivers of the permanent morphological transformations in dwarfs (e.g. Toomre, 1977; Knebe et al., 2006; Dekel et al., 2009; Naab et al., 2014; Martin et al., 2018). Furthermore, mergers are believed to be the main trigger of intense star formation in pairs (e.g. Tissera et al., 2002; Ellison et al., 2013), some of which result in the formation of starburst dwarfs, such as blue compact dwarf galaxies (BCDs; Zhang et al., 2020). Similarly, Ruiz-Lara et al. (2021) interprets the reignition of star formation after a quiescent period and the presence of metal-poor stars in Leo I as evidence of a past merger event, possibly with an ultra-faint dwarf galaxy. The dwarf-dwarf mergers can also leave discernible traces in their gas-phase metallicity. For instance, Michel-Dansac et al. (2008) showed that in a large sample of dwarf galaxy pairs with strong signs of interactions, the gas-phase metallicity is 0.2 dex higher than in their non-interacting counterparts.
@@
Observed morphological asymmetries and disturbances, such as shell-like features and tidal tails around dwarf galaxies, serve as a telltale of their relatively recent interactions (e.g. Conselice et al., 2003; Klimentowski et al., 2010; Paudel et al., 2017). For instance, Coleman et al. (2004) suggested a past merger as a possible source of the small shell-like feature around the Fornax dwarf galaxy (Yozin & Bekki, 2012). Amorisco et al. (2014) showed that the kinematically detected stellar stream in Andromeda II can be a possible remnant of a past dwarf-dwarf merger. Similarly, peculiarities in kinematics and chemical properties of other Local Group (LG) dwarf galaxies (such as Carina and Sculptor) are better explained as footprints of dwarf-dwarf encounters (Tolstoy et al., 2004; Venn et al., 2012).
@@
Detection of mergers among dwarfs extends beyond the LG (e.g. Paudel & Sengupta, 2017; Mićić et al., 2023; Román et al., 2023). For instance, Graham et al. (2012) proposed that the merger of two dwarf galaxies featuring disc-like substructures could account for the rectangular shape of LEDA074886. Paudel et al. (2017) investigated the merger origin of shell-like features in Virgo early-type dwarf galaxies. Subsequently, in Paudel et al. (2018), they shed light on the prevalence of dwarf-dwarf mergers by presenting an extensive catalogue of nearby interacting and merging dwarf galaxies up to z 0.02, but no information regarding the large-scale environment of this sample was given. Kimbro et al. (2021) highlights the 10 kpc string of massive star clusters, with stellar populations younger than 10 Myr, bridging the northern and southern components of Mrk709 and suggests that it is possibly due to the early stages of a dwarf-dwarf galaxy merger.
@@
Most of the dwarf-dwarf mergers investigated in the literature thus far are mainly located in the high-density regions (i.e. in galaxy groups and outskirts of clusters), where environmental effects are vigorous and non-negligible. In these cases, dwarf merging pairs are not only transformed due to the merging and close interaction with each other, but also experience tidal forces exerted by their host halo as well as hydrodynamical interactions with the surrounding intergalactic medium (e.g. Boselli & Gavazzi, 2014). Therefore, in such cases, disentangling between other environmental transformations and those solely induced by the merger event is not straightforward. One way to deepen our comprehension of dwarf galaxies’ growth and transformation, specifically through mergers, is to investigate dwarf-dwarf mergers in isolation (e.g. Kado-Fong et al., 2020; Grajales-Medina et al., 2023; Calderón-Castillo et al., 2024; Calderón-Castillo & Smith, 2024; Vásquez-Bustos et al., 2025) and in pristine environments, such as cosmic voids (Annibali et al., 2016). Voids represent the most under-dense regions of the cosmic web (e.g. Peebles, 2001; Kreckel et al., 2011), and they span large volumes (i.e. the average size of 35 h-1 Mpc; van de Weygaert & Platen, 2011). Within these vast expanses, galaxies are less disturbed and modified by the complex environmental processes in the galaxy groups and clusters.
@@
Few theoretical studies have focused on galaxy mergers in voids, revealing that the mean number of mergers does not depend significantly on the void environment (e.g. Rosas-Guevara et al., 2022). The only difference reported so far between mergers in voids and those in clusters or groups is that mergers in voids tend to have occurred more recently (i.e. 2 Gyr), particularly for galaxies with log(M⋆/M⊙) ¡ 9.5 (Rodríguez-Medrano et al., 2024). From observations, several studies have also reported the presence of dwarf galaxies in voids with unusual features in their morphologies, which could be indicative of past mergers (e.g. Kreckel et al., 2011). Interactions and mergers between small groups of galaxies have also been reported in some voids by a few studies (e.g. Beygu et al., 2013).
@@
In this study, we report the serendipitous detection of a gas-rich merging dwarf pair located at the central regions (i.e. at R 0.13 of void effective radius) of a cosmic void region that has been observed as part of the Calar Alto Void Integral-field Treasury Survey (CAVITY222https://cavity.caha.es/) project (Pérez et al., 2024). This legacy integral-field unit (IFU) survey aims to investigate the formation and evolution of nearby (0.005 z 0.050) void galaxies using their spatially resolved kinematics and stellar population properties (e.g. Domínguez-Gómez et al., 2023; Conrado et al., 2024). The stellar mass ratio of galaxies in this merging pair, their local and large-scale environments, and their orientation with respect to the line of sight make it a great candidate for better understanding details of merger events in low-mass galaxies. To this end, we investigate spatially resolved (on kpc scales) ionised gas kinematics, gas-phase metallicity, and SFR of this isolated merging pair and compare our results with their other counterparts in denser environments than voids. The results of this study can shed light on the crucial role of mergers in shaping the present-day properties of dwarf galaxies and better distinguish between the footprints of environmental effects and those stemming from merger events in the mass assembly of dwarf galaxies.
@@
This paper is organised as follows: In Section 2, we introduce the IFU and deep photometric data sets used for the analysis. In Section 3, we describe the methods we utilised for deriving kinematics and ionised gas emission. In Section 4, we explain the results, and in Section 5, we discuss them by comparing the interacting pair with its other counterparts from different environments in the literature. We summarise the key points of this work in Section 6. In this work, we assume a flat CDM cosmology with H0=69.6 [km s-1 Mpc-1], = 0.286, and = 0.714.
@@ 2.1 IFU data
The interacting pair was observed for a total exposure time of 5400 s divided into three dithering positions, each with two exposures of 900 s, using the PMAS/PPAK-IFU spectrograph at the 3.5 meter telescope of the Calar Alto Observatory (CAHA) as part of the CAVITY project and under the survey ID of CAVITY35844 (Pérez et al., 2024). Observations were carried out using the V500 grating of the IFU instrument, which covers the optical wavelength range from 3750 to 7500 Å with a resolving power of R 850 at 5000 Å and an average full width at half maximum (FWHM) of 6.0 Å. The instrument has a field-of-view (FOV) of 74 64 arcsec2, with a spatial sampling of 1 arcsec, that provides adequate coverage of both interacting pairs and their tidal tail (see Fig. 1). Due to the small size of the galaxy pair, it occupies only the central region of the PMAS FOV and is therefore unaffected by vignetting effects. Data reduction and construction of the final calibrated data cube were carried out through the CAVITY standard data reduction pipeline. In this process, along with other corrections and calibrations, the datacube has also been corrected for the Galactic extinction. For more details on the CAVITY data reduction pipeline, see García-Benito et al. (2024).
@@ 2.2 CAVITY+ deep optical imaging
For our analysis, we took advantage of the deep optical imaging carried out with the 2.54 m Isaac Newton Telescope (INT), as part of the CAVITY+ survey (Pérez et al., 2024). This survey uses the Wide Field Camera (WFC), mounted on the INT primary focus, that approximately covers a FOV of 34’ 34’ with a pixel scale of 0.33 arcsec pix-1. In Fig. 1 Panel A, we show CAVITY35843 and 35844 (marked with an orange box) along with CAVITY38680 (their only neighbouring galaxy within a projected search area of 1.5 Mpc, marked with a green box, see Table 1) in the reduced and calibrated INT images taken in the r-band filter. For details on the CAVITY+ deep optical imaging survey and data reduction, see Pérez et al. (2024).
@@
In Panel C of Fig. 1, we show a zoom-in view of CAVITY35843 and 35844 from DECaLS (left-hand Panel) and INT g+r images (right-hand Panel). While the data coming from INT have much higher surface brightnesses (30.0 and 29.0 mag arcsec-2 in g and r bands respectively, measured in 3 sigma, 1010 arcsec boxes; Román et al., 2020), their seeing is mediocre (1.9 and 2.0 arcsec FWHM). To address this, we utilise DECAM instrument data333Dark Energy Camera, https://noirlab.edu/science/programs/ctio/instruments/Dark-Energy-Camera, that have slightly lower surface brightness limits (approximately 1 mag arcsec-2 less Martínez-Delgado et al. 2023), but offer superior seeing conditions, with FWHM values of 1.18 and 1.29 arcsec in the r and g bands, respectively. Both data show two bright cores, representing the core of each interacting dwarf component, and a faint tidal tail heading toward the north-west. In Panel D of the same figure, we show the surface brightness map of this interacting pair using g-band INT images. The two bright cores of this interacting pair have similar surface brightness 19.5 [mag arcsec 20.5 in g-band. The tidal tail is 4 to 5 mag fainter than the two cores. Additionally, we could trace two low surface brightness (LSB) features ( 25 mag arcsec-2, marked with arrows in panel D) in the north-east and south-west of the interacting system, that are possibly the results of the ongoing interaction between these two dwarf galaxies. These LSB features are not traced in any other optical data available in the literature (including DECaLS presented here).
@@
In Panel E of Fig. 1, we show the g-r colour map of CAVITY35843 and 35844, computed using INT deep images in SDSS g and r filters. Blue regions with 0.10 g-r[mag]0.25, possibly dominated by star formation, are seen in the tidal tail and the galaxy’s east and west parts. Additionally, we trace a curved strip-like feature, dominated by red colours (g-r 0.6 mag), that separates the two bright cores of the interacting pair and their tidal tails. Compared with the DECaLS colour image in Panel C, one can notice that this red curved strip traces the distribution of dust. In the southern part of the interacting pair, we also observe a distribution of redder light. However, no dust lanes were visually detectable in these regions in the DECaLS data; thus, they may trace the presence of older or metal-richer stellar populations.
@@
In panels A and B of Fig. 2, we show the surface brightness (from the INT g-band image) and signal-to-noise ratio (S/N, extracted from PPAK data) maps of the merging pair, respectively. Here, we convolved the INT surface brightness map to match the spatial resolution of the PPAK data (i.e. with a pixel size of 1 arcsec corresponding to the spatial resolution of 2.5 arcsec). The S/N is defined on the continuum level of each spaxel’s spectrum and within the 4590 to 4610 Å wavelength range. As marked on the surface brightness map and to ease the discussion hereafter, we refer to the brightest interacting galaxy (i.e. CAVITY35843) as component A; the other (i.e. CAVITY35844) is referred to as component B. In our analysis, we excluded spaxels of the PPAK data with S/N 15 (except in the analysis in Section 4.2) to attain reliable measurements. This mainly excludes the outskirts and faint tidal tail of the merging pair.